# Practical Global and Local Bounds in Gaussian Process Regression via Chaining

### v1

@@ Abstract
Gaussian process regression (GPR) is a popular nonparametric Bayesian method that provides predictive uncertainty estimates and is widely used in safety-critical applications. While prior research has introduced various uncertainty bounds, most existing approaches require access to specific input features and rely on posterior mean and variance estimates or tuning hyperparameters. These limitations hinder robustness and fail to capture the model’s global behavior in expectation. To address these limitations, we propose a chaining-based framework for estimating upper and lower bounds on the expected extreme values over unseen data, without requiring access to specific input locations. We provide kernel-specific refinements for commonly used kernels such as RBF and Matérn, in which our bounds are tighter than generic constructions. We further improve numerical tightness by avoiding analytical relaxations. In addition to global estimation, we also develop a novel method for local uncertainty quantification at specified inputs. This approach leverages chaining geometry through partition diameters, adapting to local structure without relying on posterior variance scaling. Our experimental results validate the theoretical findings and demonstrate that our method outperforms existing approaches on both synthetic and real-world datasets.
@@ Introduction
Despite this progress, most existing approaches aim to control errors or construct uncertainty intervals at specified input locations. These bounds typically require access to specific test input features and rely on posterior mean and variance or intensive hyperparameter tuning, limiting their adaptability and often resulting in conservative or unstable intervals. Moreover, they do not directly address global behavior, such as the expected maximum bound, which is critical for assessing model reliability in safety-critical settings. For instance, autonomous systems must ensure trajectory deviations remain within safety thresholds over time, while disaster and financial risk models focus on expected peaks, such as the highest flood level or the worst market loss.
To address these limitations, we propose a novel framework that directly controls the global behavior of Gaussian processes without requiring access to specific input locations. Specifically, we derive an expected upper bound on the global maximum of the process, which, by symmetry, also provides a corresponding lower bound. To the best of our knowledge, this is the first application of chaining-based bounds in the context of GPR.
In practical applications, loose bounds can lead to inefficient decisions and resource waste, motivating the need for tighter estimates. We further refine the chaining bounds by explicitly leveraging the properties of kernels such as RBF and Matérn, resulting in provably tighter and more practical bounds. Our implementation avoids analytical relaxations such as loose constant factors that are typically introduced for mathematical convenience but often result in overly conservative estimates.
While the above method focuses on global behavior, we also propose a complementary approach to quantify uncertainty at specific input locations. Instead of relying on chaining accumulation, this method leverages geometric partitioning inspired by chaining to define locally adaptive reference sets. The uncertainty bounds are then computed by scaling the diameters of these partitions, aligning with the local geometry of the input space and avoiding direct reliance on posterior mean and variance scaling.
@@
To illustrate, consider a stochastic process $(X_{t})_{t\in T}$, in which the difference between $X_{t}$ and $X_{0}$ is expressed as $X_{t}-X_{0}=\sum_{n\geq 1}\left(X_{t}-X_{t-1}\right)$. When many variables $X_{t}$ in $T$ are nearly identical, strong correlations between them can obscure the true variation in the process. Grouping similar variables together helps reduce this redundancy by allowing us to approximate these highly correlated variables with a representative value, thereby simplifying the analysis and making the process easier to interpret and work with. A more detailed explanation can be found in Appendix A.1.
For $n\geq 0$, we select a subset $T_{n}$, and for each $t\in T$, we choose an approximation $\pi_{n}(t)$ from $T_{n}$. Using these $\pi_{n}(t)$ points, we obtain the corresponding $X_{\pi_{n}(t)}$ variables, which serve as successive approximations of $X_{t}$. We start by assuming that $T_{0}$ contains only one element $t_{0}$, and thus $\pi_{0}(t)=t_{0}$ for all $t\in T$. The core relation is: $X_{t}-X_{t_{0}}=\sum_{n\geq 1}\left(X_{\pi_{n}(t)}-X_{\pi_{n-1}(t)}\right).$ This equality holds because, for sufficiently large $n$, $\pi_{n}(t)$ equals $t$, meaning that beyond a certain point, the approximation stops, and the series becomes a finite sum. Specifically, as $n$ increases, the sets $T_{n}$ become progressively finer, eventually covering all points in $T$. Once $T_{n}$ contains $t$, we have $\pi_{n}(t)=t$, so no new information is added by further terms in the series. As a result, the infinite series truncates to a finite sum. This ensures convergence in practical settings where the process $X_{t}$ is fully captured after a finite number of terms. The efficacy of this approach is rooted in the fact that for each approximation $\pi_{i}(t)$, the variables $X_{t}-X_{\pi_{i}(t)}$ are smaller than $X_{t}-X_{t_{0}}$, making their supremum easier to handle. We present a simple example in Appendix A.2.
@@
The concept of bounds in GPR originates from the Bayesian confidence intervals, which assume the target function is drawn from a Gaussian process prior, reflect posterior uncertainty, and are widely used in traditional GPR setups.
In contrast, the frequentist literature develops high-probability uncertainty bounds, which do not assume a prior but instead provide guarantees over repeated sampling. For example, srinivas2010gaussian, srinivas2012information, and chowdhury2017kernelized derive time-uniform bounds on the prediction error by leveraging information gain and RKHS norms. These bounds are often used in Bayesian optimization and bandit settings. Similarly, fiedler2021practical refine these bounds by introducing a model-misspecification-aware error term, still within a frequentist framework. A different class of results involves deterministic error bounds from the field of scattered data approximation. Classical works such as wu1993local and schaback1999improved, as summarized in wendland2004scattered, provide interpolation-based error bounds using fill distance and smoothness assumptions. These results are fundamentally non-probabilistic and rely on properties of the function space (e.g., RKHS regularity), assuming noiseless observations and no distributional randomness.
@@
These methods bound uncertainty pointwise but not global deviations. Our work advances GPR bounds in two directions. First, we implement Talagrand’s generic chaining bound to estimate the global expected supremum error, capturing expected global deviation rather than input-specific guarantees. While existing frequentist bounds upper-bound this supremum by maximizing over uncertainty intervals at given inputs, our method offers a theoretical alternative that targets the expected supremum. Second, we propose a new method for uncertainty quantification at specified inputs, inspired by chaining geometry, which uses partitions to construct locally adaptive bounds around each input, reducing reliance on global posterior variance. These methods jointly extend GPR bounds from local to global uncertainty control.
@@
We now present our technical contributions. We avoid posterior variance scaling compare to existing methods and apply chaining directly to the distance metric induced by the kernel function. While GPR typically fits data and selects kernel hyperparameters, we use the kernel solely to define a distance metric, bypassing posterior inference. This eliminates dependence on posterior variance and calibration, yielding tighter, more reliable bounds. We assume a bounded RKHS function and i.i.d. sub-Gaussian noise.
@@
(Generic Bound) Theorem 1, combined with the formula $\mathbb{E}\left[Y\right]=\int_{0}^{\infty}P(Y\geq u)\,du,$ which gives the expectation of a non-negative random variable, leads to the following upper bound:
@@
Proof.
(Sketch) Theorem 1 combined with the integral formula for non-negative expectations to get the upper bound $\mathbb{E}[\sup_{t\in T}|X_{t}-X_{t_{0}}|]\leq LS\sqrt{\pi/2}$. The approximation $d(t,\pi_{n}(t))=\inf_{s\in T_{n}}d(t,s)$ and triangle inequality control chaining increments. Detailed proofs of Theorem 1 and Theorem 2 are in Appendix B.1 and Appendix B.2. ∎
@@
When $\nu=p+1/2$ with $p\in\mathbb{N}$, the kernel simplifies to a product of an exponential and a polynomial of order $p$ (seeger2004gaussian). A commonly used case is $\nu=3/2$, where
@@
The distance between two points $s$ and $t$ in the context of GPs is defined as $d(s,t)=\sqrt{\mathbb{E}[(X_{s}-X_{t})^{2}]}$, where $X_{s}$ and $X_{t}$ are the values at points $s$ and $t$ respectively. This distance metric is derived from the covariance function $K(s,t)$, which describes the covariance between the random variables $X_{s}$ and $X_{t}$. Specifically, it can be expanded as:
@@
Proof.
(Sketch) We used the kernel-induced distance as $d(s,t)^{2}=2\sigma^{2}(1-\exp(-\|s-t\|^{2}/2l^{2}))$, and apply the inequality $\|s-t\|^{2}+\|t-u\|^{2}\geq\|s-u\|^{2}/2$ with the convexity of the exponential to get $d(s,u)^{2}\leq d^{\prime}(s,t)^{2}+d^{\prime}(t,u)^{2}$, where $d^{\prime}(s,t)^{2}=2\sigma^{2}-2\sigma K^{1/2}(s,t)$. We use the approximation and triangle inequality as in Theorem 2. Detailed proof is given in Appendix B.3. ∎
While the RBF kernel is widely used, other kernels, such as the Matérn kernel, are better suited for specific applications. In the following, we provide and prove the upper and lower bounds for the Matérn kernel with $\nu=3/2$.
Lemma 4.
(Tighter Matérn Bound) Consider a Gaussian process $(X_{t})_{t\in T}$ with a Matérn kernel $K(s,t)=\left(1+\frac{\sqrt{3}\|s-t\|}{l}\right)\exp\left(-\frac{\sqrt{3}\|s-t\|}{l}\right)$ where $T$ is an input/index set, $\|s-t\|$ is the Euclidean distance between input points $s\in T$ and $t\in T$, and $l$ is the length-scale parameter. Then
@@
Proof.
(Sketch) By applying Chebyshev’s sum inequality to the sequences $(1+x_{i})$ and $\exp(-x_{i})$ with $x_{i}=\frac{\sqrt{3}\|\cdot\|}{l}$, and leveraging the monotonicity of the function $f(x)=(1+x)\exp(-x)$, we establish that the kernel satisfies a relaxed subadditivity condition $d(s,u)^{2}\leq d^{\prime}(s,t)^{2}+d^{\prime}(t,u)^{2}-2,$ where $d^{\prime}(s,t)^{2}=2-2K^{\prime}(s,t)$. We then follow the approximation and triangle inequality argument outlined in Theorem 2. Detailed proof is provided in Appendix B.4. ∎
@@
However, beyond kernel-specific flexibility, we also address a more fundamental limitation in classical chaining theory. Generic chaining is not optimized for sharp constants, resulting in conservative numerical bounds. In particular, classical results (e.g., proof in Appendix B.1) often include a large prefactor $L=\exp(9/2)$.
@@
where $S\leq(1+\sqrt{2})\sup_{t\in T}\sum_{n\geq 0}2^{n/2}d^{\prime}(t,s)$ for RBF kernel in Theorem 3; $S\leq(1+\sqrt{2})\sup_{t\in T}\sum_{n\geq 0}2^{n/2}d^{\prime\prime}(t,s)$ for Matérn kernel in Theorem 4.
Proof.
(Sketch) The Gaussian tail bound is applied to the increment $X_{\pi_{n}(t)}-X_{\pi_{n-1}(t)}$, bounding its tail probability by $\exp(-u^{2}2^{n-1})$, where the threshold is $u\,2^{n/2}d(\pi_{n}(t),\pi_{n-1}(t))$. A union bound $p(u)$ on the failure probability of the event $\Omega_{u}^{c}$. To avoid introducing large analytic prefactors $L$, noting that $p(u)\leq 1$, so we can get $\mathbb{E}[\sup_{t\in T}|X_{t}-X_{t_{0}}|]\leq\int_{0}^{\infty}\min(p(u/S),1)\,du,$, where $S$ is bounded via Lemmas 3 and 4. Detailed proof is provided in Appendix B.5. ∎
@@
When incorporating test set features to predict pointwise bounds, the objective shifts from inferring the global supremum (and infimum) to estimating bounds for specific test points, $t^{\prime}$. Symmetric intervals around a fixed reference (e.g., $X_{t_{0}}\approx 0$) may lack tightness, as the uncertainty should instead be symmetric around $X_{t^{\prime}}$. To address this, we focus on $X_{t}-X_{\pi_{n}(t)}$, where $X_{\pi_{n}(t)}$ is a reference point of $X_{t}$ in $T_{n}$, yielding tighter and more precise bounds. Such bounds are useful when we wish to bound the uncertainty around the predicted value of an individual test point.
@@
Proof.
(Sketch) By introducing the confidence parameter $\delta=2\exp(-u^{2}2^{n-1})$ and taking the natural logarithm of both sides, we have $\ln(\delta)=\ln(2)-u^{2}2^{n-1}.$ Rearranging for $u^{2}$, we get $u=\sqrt{\frac{\ln(2)+\ln(1/\delta)}{2^{n-1}}}.$ and substituting this expression for $u$ back into the inequality for $|X_{t}-X_{\pi_{n}(t)}|$ gives the result. Detailed proof is in Appendix B.6. ∎
In the chaining method, it is challenging to construct the sets $T_{n}$. talagrand2014upper reformulate this as a geometric problem, replacing $d(t,T_{n})$ with partition diameters. This approach replaces the inherently combinatorial nature of supremum computations with geometric quantities that are easier to control and analyze. Partition diameters, as global quantities, reduce computational complexity to a manageable recursive process and provide better theoretical control.
To formalize this approach, it define an admissible sequence $(A_{n})_{n\geq 0}$ as an increasing sequence of partitions of $T$, where $A_{0}=\{T\}$, and for $n\geq 1$, $|(A_{n})|\leq 2^{2^{n}}$. Each $A_{n}(t)$ denotes the unique set in $A_{n}$ containing $t\in T$.
From each partition $A_{n}$, a representative point is selected from every set $A\in A_{n}$ to construct the subset $T_{n}\subseteq T$. By construction, for any $t\in T$ and $n\geq 0$, the distance between $t$ and $T_{n}$ satisfies $d(t,T_{n})\leq\Delta(A_{n}(t)),$ where $\Delta(A_{n}(t))$ represents the diameter of $A_{n}(t)$ under the metric $d$. So Theorem 6 can be expressed as:
@@
17:  Compute uncertainty bound $u(t)=\Delta(A_{n_{\text{max}},k^{*}})\cdot\sqrt{2\cdot(\log(1/\delta)+\log(2))}$ using Theorem 6.
@@
We compare our chaining method to the following three state-of-the-art baselines: (i) Lederer19 (lederer2019uniform), which introduces probabilistic Lipschitz constants to reduce the reliance on prior knowledge, estimates errors on a finite grid, and extends them to the input space; (ii) Fiedler21 (fiedler2021practical), which modifies its objective bound function by introducing an error term based on the work of (chowdhury2017kernelized); (iii) Capone22 (capone2022gaussian), which tackles hyperparameter misspecification by proposing a method to calculate error bounds across a given range of hyperparameters; and (iv) Bayesian CI, which uses the GP posterior mean and standard deviation to form Bayesian credible intervals $\mu(x)\pm z\cdot\sigma(x)$ .
@@
For Lederer19, we use the default implementation with 10000 grid points over a fixed domain and resolution $\tau=10^{-8}$. For Fiedler21, we compare the default RKHS norm $B=2$ with a data-driven estimate $y^{\top}(K+\lambda I)^{-1}y$, using the better one, and fix the noise level to $0.0001$. For Capone22, we use provided hyperparameter grids when available, or adopt ranges from similar datasets otherwise. All methods use the RBF kernel as in their original versions.

### v2

@@ Abstract
Gaussian process regression (GPR) is a popular nonparametric Bayesian method that provides predictive uncertainty estimates and is widely used in safety-critical applications. While prior research has introduced various uncertainty bounds, most existing approaches require access to specific input features, and rely on posterior mean and variance estimates or the tuning of hyperparameters. These limitations hinder robustness and fail to capture the model’s global behavior in expectation. To address these limitations, we propose a chaining-based framework for estimating upper and lower bounds on the expected extreme values over unseen data, without requiring access to specific input features. We provide kernel-specific refinements for commonly used kernels such as RBF and Matérn, in which our bounds are tighter than generic constructions. We further improve numerical tightness by avoiding analytical relaxations. In addition to global estimation, we also develop a novel method for local uncertainty quantification at specified inputs. This approach leverages chaining geometry through partition diameters, adapting to local structures without relying on posterior variance scaling. Our experimental results validate the theoretical findings and demonstrate that our method outperforms existing approaches on both synthetic and real-world datasets.
@@ Introduction
Despite this progress, most existing approaches aim to control errors or construct uncertainty intervals at specified input locations. These bounds typically require access to specific test input features, and rely on posterior mean and variance or intensive hyperparameter tuning, limiting their adaptability and often resulting in conservative or unstable intervals. Moreover, they do not directly address global behavior, such as the expected maximum bound, which is critical for assessing model reliability in safety-critical settings. For instance, autonomous systems must ensure trajectory deviations remain within safety thresholds over time, while disaster and financial risk models focus on expected peaks, such as the highest flood level or the worst market loss.
To address these limitations, we propose a novel framework that directly controls the global behavior of Gaussian processes without requiring access to specific input features. Specifically, we derive an expected upper bound on the global maximum of the process, which, by symmetry, also provides a corresponding lower bound. To our knowledge, this is the first application of chaining-based bounds in the context of GPR.
In practical applications, loose bounds can lead to inefficient decisions and resource waste, motivating the need for tighter estimates. We tighten the chaining bounds by exploiting properties of kernels, such as RBF and Matérn, yielding provably tighter and more practical bounds. Our implementation avoids analytical relaxations, such as loose constant factors, that are typically introduced for mathematical convenience but often result in overly conservative estimates.
While our method captures global behavior, we also propose a complementary method to quantify uncertainty at specific input features. Instead of relying on chaining accumulation, this method is inspired by chaining and leverages geometric partitioning to define locally adaptive reference sets. The uncertainty bounds are then computed by scaling the diameters of these partitions, aligning with the local geometry of the input space, and avoiding direct reliance on posterior mean and variance scaling.
@@
To illustrate, consider a stochastic process $(X_{t})_{t\in T}$, in which the difference between $X_{n}$ and $X_{0}$ is expressed as $X_{n}-X_{0}=\sum_{t=1}^{n}\left(X_{t}-X_{t-1}\right)$. When many variables $X_{t}$ in $T$ are nearly identical, strong correlations between them can obscure the true variation in the process. Grouping similar variables together helps reduce this redundancy by allowing us to approximate these highly correlated variables with a representative value, thereby simplifying the analysis and making the process easier to interpret and work with. A more detailed explanation can be found in Appendix A.1.
For $n\geq 0$, we select a subset $T_{n}$, and for each $t\in T$, we choose an approximation $\pi_{n}(t)$ from $T_{n}$. Using these $\{\pi_{n}(t)\}$ points, we obtain the corresponding $\{X_{\pi_{n}(t)}\}$ variables, which serve as successive approximations of $\{X_{t}\}$. We start by assuming that $T_{0}$ contains only one element $t_{0}$, and thus $\pi_{0}(t)=t_{0}$ for all $t\in T$. The core relation is: $X_{t}-X_{t_{0}}=\sum_{n\geq 1}\left(X_{\pi_{n}(t)}-X_{\pi_{n-1}(t)}\right).$
This equality holds because, for sufficiently large $n$, $\pi_{n}(t)$ equals $t$, meaning that beyond a certain point, the approximation stops, and the series becomes a finite sum. Specifically, as $n$ increases, the sets $T_{n}$ become progressively finer, eventually covering all points in $T$. Once $T_{n}$ contains $t$, we have $\pi_{n}(t)=t$, so no new information is added by further terms in the series. As a result, the infinite series truncates to a finite sum. This ensures convergence in practical settings where the process $X_{t}$ is fully captured after a finite number of terms. The efficacy of this approach is rooted in the fact that for each approximation $\pi_{i}(t)$, the variables $X_{t}-X_{\pi_{i}(t)}$ are smaller than $X_{t}-X_{t_{0}}$, making their supremum easier to handle. We present a simple example in Appendix A.2.
@@
The concept of bounds in GPR originates from the Bayesian confidence intervals, which assume the target function is drawn from a Gaussian process prior and are widely used in traditional GPR setups to reflect posterior uncertainty.
In contrast, the frequentist literature develops high-probability uncertainty bounds, which do not assume a prior but instead provide guarantees over repeated sampling. For example, srinivas2010gaussian, srinivas2012information, and chowdhury2017kernelized derive time-uniform bounds on the prediction error by leveraging information gain and RKHS norms. These bounds are often used in Bayesian optimization and bandit settings. Similarly, fiedler2021practical refine these bounds by introducing a model-misspecification-aware error term, still within a frequentist framework. A different class of results involves deterministic error bounds from the field of scattered data approximation. Classical works such as wu1993local and schaback1999improved, as summarized in wendland2004scattered, provide interpolation-based error bounds using fill distance and smoothness assumptions. These results are fundamentally non-probabilistic, rely on properties of the function space (e.g., RKHS regularity), and assume noiseless observations and the absence of distributional randomness.
@@
These methods bound uncertainty pointwise but not global deviations. Our work advances GPR bounds in two directions. First, we implement Talagrand’s generic chaining bound to estimate the global expected supremum error, capturing expected global deviation rather than input-specific guarantees. While existing frequentist bounds upper-bound this supremum by maximizing over uncertainty intervals at given inputs, our method offers a theoretical alternative that targets the expected supremum. Second, we propose a new method for uncertainty quantification at specified inputs; Inspired by chaining geometry, it uses partitions to construct locally adaptive bounds around each input, reducing reliance on global posterior variance. Our methods jointly extend GPR bounds from local to global uncertainty control.
@@
We now present our technical contributions. We avoid posterior variance scaling compared to existing methods and apply chaining directly to the distance metric induced by the kernel function. While GPR typically fits data and selects kernel hyperparameters, we use the kernel solely to define a distance metric, bypassing posterior inference. This eliminates dependence on posterior variance and calibration, yielding tighter, more reliable bounds. We assume a bounded RKHS function and i.i.d. sub-Gaussian noise.
@@
(Talagrand’s Generic Bound) Theorem 1, combined with the formula $\mathbb{E}\left[Y\right]=\int_{0}^{\infty}P(Y\geq u)\,du,$ which gives the expectation of a non-negative random variable, leads to the following upper bound:
@@
Proof sketch..
We combine theorem 1 with the integral formula for non-negative expectations to get the upper bound $\mathbb{E}[\sup_{t\in T}|X_{t}-X_{t_{0}}|]\leq LS\sqrt{\pi/2}$. The approximation $d(t,\pi_{n}(t))=\inf_{s\in T_{n}}d(t,s)$ and the triangle inequality control chaining increments. Detailed proofs of Theorems  1 and  2 are in Appendices  B.1 and  B.2. ∎
@@
When $\nu=p+1/2$ with $p\in\mathbb{N}$, the kernel simplifies to a product of an exponential and a polynomial of order $p$ (seeger2004gaussian). Commonly $\nu=3/2$, giving
@@
The distance between two points $s$ and $t$ in the context of GPs is $d(s,t)=\sqrt{\mathbb{E}[(X_{s}-X_{t})^{2}]}$, where $X_{s}$ and $X_{t}$ are the values at points $s$ and $t$ respectively. This distance metric is derived from the covariance function $K(s,t)$, which describes the covariance between the random variables $X_{s}$ and $X_{t}$. Specifically, it can be expanded as:
@@
Proof sketch..
We use the kernel-induced distance $d(s,t)^{2}=2\sigma^{2}(1-\exp(-\|s-t\|^{2}/2l^{2}))$, and apply the inequality $\|s-t\|^{2}+\|t-u\|^{2}\geq\|s-u\|^{2}/2$ with the convexity of the exponential to get $d(s,u)^{2}\leq d^{\prime}(s,t)^{2}+d^{\prime}(t,u)^{2}$, where $d^{\prime}(s,t)^{2}=2\sigma^{2}-2\sigma K^{1/2}(s,t)$. We use the approximation and triangle inequality. A detailed proof is given in Appendix B.3. ∎
While the RBF kernel is widely used, other kernels, such as the Matérn kernel, are better suited for specific applications. In the following, we prove the upper and lower bounds for the Matérn kernel with $\nu=3/2$.
Theorem 4.
(Tighter Matérn Bound) Consider a Gaussian process $(X_{t})_{t\in T}$ with a Matérn kernel $K(s,t)=\left(1+\frac{\sqrt{3}\|s-t\|}{l}\right)\exp\left(-\frac{\sqrt{3}\|s-t\|}{l}\right)$, where $T$ is an input/index set, $\|s-t\|$ is the Euclidean distance between input points $s\in T$ and $t\in T$, and $l$ is the length-scale parameter. Then
@@
Proof sketch..
By applying Chebyshev’s sum inequality to the sequences $(1+x_{i})$ and $\exp(-x_{i})$ with $x_{i}=\frac{\sqrt{3}\|\cdot\|}{l}$, and leveraging the monotonicity of the function $f(x)=(1+x)\exp(-x)$, we establish that the kernel satisfies a relaxed subadditivity condition $d(s,u)^{2}\leq d^{\prime}(s,t)^{2}+d^{\prime}(t,u)^{2}-2,$ where $d^{\prime}(s,t)^{2}=2-2K^{\prime}(s,t)$. We then follow the approximation and triangle inequality. A detailed proof is provided in Appendix B.4. ∎
@@
However, beyond kernel-specific flexibility, we also address a more fundamental limitation in classical chaining theory. Generic chaining is not optimized for sharp constants, resulting in loose conservative numerical bounds. In particular, classical results (e.g., the proof in Appendix B.1) often include a large prefactor $L=\exp(9/2)$.
@@
where $S\leq(1+\sqrt{2})\sup_{t\in T}\sum_{n\geq 0}2^{n/2}d^{\prime}(t,s)$ for the RBF kernel in Theorem 3; $S\leq(1+\sqrt{2})\sup_{t\in T}\sum_{n\geq 0}2^{n/2}d^{\prime\prime}(t,s)$ for the Matérn kernel in Theorem 4.
Proof sketch..
The Gaussian tail bound is applied to the increment $X_{\pi_{n}(t)}-X_{\pi_{n-1}(t)}$, bounding its tail probability by $\exp(-u^{2}2^{n-1})$, where the threshold is $u\,2^{n/2}d(\pi_{n}(t),\pi_{n-1}(t))$. A union bound $p(u)$ is used for the failure probability of the event $\Omega_{u}^{c}$. To avoid introducing large analytic prefactors $L$. We note that $p(u)\leq 1$ to get $\mathbb{E}[\sup_{t\in T}|X_{t}-X_{t_{0}}|]\leq\int_{0}^{\infty}\min(p(u/S),1)\,du,$, where $S$ is bounded via theorems 3 and 4. A detailed proof is provided in Appendix B.5. ∎
@@
When incorporating test set features to predict pointwise bounds, the objective shifts from inferring the global supremum (and infimum) to estimating bounds for a specific test point, $t^{\prime}$. Symmetric intervals around a fixed reference (e.g., $X_{t_{0}}\approx 0$) may lack tightness, as the uncertainty should instead be symmetric around $X_{t^{\prime}}$. To address this, we focus on $X_{t}-X_{\pi_{n}(t)}$, where $X_{\pi_{n}(t)}$ is a reference point of $X_{t}$ in $T_{n}$, yielding tighter and more precise bounds. Such bounds are useful when we wish to bound the uncertainty around the predicted value of an individual test point.
@@
Proof sketch..
By introducing the confidence parameter $\delta=2\exp(-u^{2}2^{n-1})$ and taking the natural logarithm of both sides, we have $\ln(\delta)=\ln(2)-u^{2}2^{n-1}.$ Rearranging for $u^{2}$, we get $u=\sqrt{\frac{\ln(2)+\ln(1/\delta)}{2^{n-1}}}.$ and substituting this expression for $u$ into the inequality for $|X_{t}-X_{\pi_{n}(t)}|$ gives the result. A detailed proof is in Appendix B.6. ∎
In the chaining method, it is challenging to construct the sets $T_{n}$. talagrand2014upper reformulates this as a geometric problem, replacing $d(t,T_{n})$ with partition diameters. This approach replaces the inherently combinatorial nature of supremum computations with geometric quantities that are easier to control and analyze. Partition diameters, as global quantities, reduce computational complexity to a manageable recursive process and provide better theoretical control.
To formalize this approach, an admissible sequence $(A_{n})_{n\geq 0}$ is defined as an increasing sequence of partitions of $T$, where $A_{0}=\{T\}$, and $|(A_{n})|\leq 2^{2^{n}}$ for $n\geq 1$. Each $A_{n}(t)$ denotes the unique set in $A_{n}$ containing $t\in T$.
From each partition $A_{n}$, a representative point is selected from every set $A\in A_{n}$ to construct the subset $T_{n}\subseteq T$. By construction, for any $t\in T$ and $n\geq 0$, the distance between $t$ and $T_{n}$ satisfies $d(t,T_{n})\leq\Delta(A_{n}(t)),$ where $\Delta(A_{n}(t))$ represents the diameter of $A_{n}(t)$ under the metric $d$. Thus Theorem 6 can be expressed as:
@@
17:  Compute uncertainty bound $u(t)=\Delta(A_{n_{\text{max}},k^{*}})\cdot\sqrt{2\cdot(\log(1/\delta)+\log(2))}$.
@@
We compare our chaining method to the following three state-of-the-art baselines, previously introduced in the related-work section: (i) Lederer19 (lederer2019uniform), which introduces probabilistic Lipschitz constants to reduce the reliance on prior knowledge, estimates errors on a finite grid, and extends them to the input space; (ii) Fiedler21 (fiedler2021practical), which modifies its objective bound function by introducing an error term based on the work of (chowdhury2017kernelized); (iii) Capone22 (capone2022gaussian), which tackles hyperparameter misspecification by proposing a method to calculate error bounds across a given range of hyperparameters; and (iv) Bayesian CI, which uses the standard GP posterior mean and standard deviation to form Bayesian credible intervals $\mu(x)\pm z\cdot\sigma(x)$.
@@
For Lederer19, we use the default implementation with 10000 grid points over a fixed domain and resolution $\tau=10^{-8}$. For Fiedler21, we compare the default RKHS norm $B=2$ with a data-driven estimate $y^{\top}(K+\lambda I)^{-1}y$, using the better one, and fix the noise level to $0.0001$. For Capone22, we use the provided hyperparameter grids when available, or adopt ranges from similar datasets otherwise. All methods use the RBF kernel as in their original versions.